OpenStudy (anonymous):

FAN AND MEDAL!!! In a direct variation, y is 9 when x is eighteen. What is y when x equals six?

OpenStudy (anonymous):

y is 3 here...

OpenStudy (imstuck):

Do you know how to set up these types of problems? with a K?

OpenStudy (anonymous):

nope

OpenStudy (imstuck):

Giving an answer doesn't help the learning process.

OpenStudy (anonymous):

yea bro

OpenStudy (imstuck):

Ok, in a direct variation problem the set up is y = kx. They tell you that y = 9 when x = 18. So let's fill that in: 9 = 18k. Solving for k we get 1/2. K = 1/2. Now use that value of k in the next part of the problem.

OpenStudy (imstuck):

What is y when x = 6? Going back to our variation equation y = kx we need to solve for y when x = 6 and we have solved for k. Solve y = kx when k = 1/2 and x = 6. \[y=\frac{ 1 }{ 2 }(6)\]What is one-half of 6?